# Signal Recovery From Product of Two Vandermonde Matrices

### 27 Jan 2024 v1

@@ Abstract
In this work, we present some new results for compressed sensing and phase retrieval. For compressed sensing, it is shown that if the unknown $n$-dimensional vector can be expressed as a linear combination of $s$ unknown Vandermonde vectors (with Fourier vectors as a special case) and the measurement matrix is a Vandermonde matrix, exact recovery of the vector with $2s$ measurements and $O(\mathrm{poly}(s))$ complexity is possible when $n\geq 2s$. Hence, in this case, there is no need for random measurements (with the number of measurements scaled by $\log(n)$) as required by conventional compressed sensing theory. From these results, a measurement matrix is constructed from which it is possible to recover $s$-sparse $n$-dimensional vectors for $n\geq 2s$ with as few as $2s$ measurements and with a recovery algorithm of $O(\mathrm{poly}(s))$ complexity. This seems to be the first measurement matrix construction achieving exact recovery of $s$-sparse vectors with $O(s)$ measurements (achieving the lower bound of $2s$) and $O(\mathrm{poly}(s))$ complexity (typically, the recovery complexity is $O(\mathrm{poly}(n))$), significantly improving upon a classical construction by Bourgain et al. In the second part of the work, these results are extended to the challenging problem of phase retrieval (compressed sensing with magnitude-only observations). The most significant discovery in this direction is that if the unknown $n$-dimensional vector is composed of $s$ frequencies with at least one being non-harmonic, $n\geq 4s-1$ and we take at least $38s-$ 3 Fourier measurements, there are, remarkably, only two possible vectors producing the observed measurement values and they are easily obtainable from each other. The two vectors can be found by an algorithm with only $O(\mathrm{poly}(s))$ complexity. This result is interesting when put in relation to recent work by Xu which shows that solving Fourier phase retrieval with $O(\mathrm{poly}(n))$ complexity proves that P = NP holds for the well known P vs NP problem. An immediate application of the new result is construction of a measurement matrix from which it is possible to recover all $s$-sparse $n$-dimensional signals (up to a global phase) from $O(s)$ magnitude-only measurements and $O(\mathrm{poly}(s))$ recovery complexity when $n\geq 4s-1$. As a consequence, the proposed matrix construction solves (and significantly improves upon) an open problem by Xu on finding a measurement matrix construction from which one can (provably) recover any $s$-sparse vector with $O(s\log(n/s))$ phase-less measurements and 1-norm minimization.
@@ I Introduction
where $\mathbb{C}^{n}_{\leq s}$ is the space of $n$-dimensional complex vectors with at most $s$ elements being non-zero and $\bm{A}$ is a known $m\times n$ measurement matrix ($m$ is the number of measurements). This problem has a rich history. Seminal works [1, 2, 3] and references therein show that, remarkably, there exist matrices $\bm{A}$ such that $\bm{x}$ can be recovered with far less measurements than $n$. We will refer to exact recovery of $\bm{x}$ as recovery with probability 1 (abbreviated as w.p.1) while probabilistic recovery refers to recovery at some probability strictly smaller than 1. The obvious 0-norm minimization that searches across all subsets of $s$ columns from $\bm{A}$ and inverts them to obtain a candidate solution is of exponential complexity. Fortunately, the work in [4] shows that there exists a recovery algorithm for (1) with $O(\mathrm{poly}(n))$ complexity, where $O(.)$ stands for big O notation. More specifically, [4] shows that, for a given $\bm{y}=\bm{A}\bm{x}$, the unique solution to the convex 1-norm optimization problem
@@
In many practical instances of compressed sensing and phase retrieval, $\bm{A}$ is constrained to be a Fourier matrix [1, 17, 11] which in practice corresponds to an apparatus that takes frequency measurements of an underlying signal. Compressed sensing and (sparse) phase retrieval with Fourier measurement matrices, known as Fourier compressed sensing and (sparse) Fourier phase retrieval, respectively, are notoriously difficult problems, as described in [1, 6, 17]. The main challenge is in obtaining an affordable recovery complexity (e.g. $O(\mathrm{poly}(n))$) with few measurements (e.g. $O(s\log(n/s))$). The breakthrough work in compressed sensing [1] considered random Fourier measurements in order to escape the deterministic nature of the problem and instead provide probabilistic recovery with complexity $O(\mathrm{poly}(n))$ and $O(s\log(n/s))$ measurements. The contributions in this work on compressed sensing show that if $\bm{x}$ is instead $s$-sparse in a Vandermonde basis, with $n\geq 2s$, then it can be recovered exactly from $O(s)$ deterministic Fourier measurements and $O(\mathrm{poly}(s))$ recovery complexity; hence, there is no need for random measurements in this case. In case of (non-sparse) Fourier phase retrieval, as shown in [6], the reason for why no exact $O(\mathrm{poly}(n))$ complexity algorithms have been proposed so far is because in doing so one solves the well known P vs NP problem, proving that P = NP holds. Interestingly, the contributions in this work on Fourier phase retrieval show that if $\bm{x}$ is $s$-sparse in a non-DFT basis, with $n\geq 4s-1$, then it can be recovered from $O(s)$ measurements and $O(\mathrm{poly}(s))$ recovery complexity.
@@ I-A Contributions
Let $\mathbb{T}=\{z;z\in\mathbb{C},|z|=1\}$ be the set of all points on the complex unit circle and (.$(.)^{T}$ the matrix transpose operator. When $z_{k}\in\mathbb{T}$, $10\leq k\leq m-$ 1, the Vandermonde matrix becomes a Fourier matrix, which we for clarity henceforth denote as $\bm{F}(\bm{z})$ (with DFT matrix as a special case when $z_{k}$ are $n$th roots of unity and $m=n$). For Vandermonde and Fourier measurement matrices, we have the following main contributions
@@
      For arbitrary $\bm{z}\in\mathbb{C}^{m}$, $m\geq 3s$ suffices.
@@
      If the elements in $\bm{\theta}$ are $n$th roots of unity ($\bm{F}(\bm{\theta})$ equals a subset of $s$ columns from a $n\times n$ DFT matrix), we have exact recovery of $\bm{\theta}$ and $|\bm{g}|$ (up to a global real scalar) from $\bm{y}$ and $\bm{A}$ with $O(\mathrm{poly}(s))$ recovery complexity. There are $2^{s-1}$ different $\bm{g}$ (up to a global phase), all having the same value of $|\bm{g}|$, producing the same $\bm{y}$.
@@
      If not all elements in $\bm{\theta}$ are $n$th roots of unity then
@@
        When the elements in $\bm{z}$ are $n$th roots of unity and $m=4s-1$, we have exact recovery of $\bm{\theta}$ and $|\bm{g}|$ (up to a global real scalar) from $\bm{y}$ and $\bm{A}$ with $O(\mathrm{poly}(s))$ recovery complexity. There are $2^{s-1}$ different $\bm{g}$ (up to a global phase), all having the same value of $|\bm{g}|$, producing the same $\bm{y}$.
@@
        For arbitrary $\bm{z}\in\mathbb{T}^{m}$ with $m\geq 8s-3$, we have exact recovery of $\bm{\theta}$ and $|\bm{g}|$ (up to a global real scalar) from $\bm{y}$ and $\bm{A}$ with $O(\mathrm{poly}(s))$ recovery complexity. There are only two different $\bm{g}$ (up to a global phase), both having the same value of $|\bm{g}|$, that produce the same $\bm{y}$; moreover, the two solutions are easily obtainable from each other.
By an arbitrary vector $\bm{a}\in\mathbb{C}^{k}$, we mean a vector chosen at random with some probability distribution that is continuous on $\mathbb{C}^{k}$.
@@
For phase-less measurements, the result in 2. states that $\bm{\theta}$ and $c|\bm{g}|^{2}$, for unknown real scalar $c$, can be found with $O(s)$ measurements and $O(\mathrm{poly}(s))$ recovery complexity. Thus, in scenarios where only the frequency components $\bm{\theta}$ and their relative magnitudes are sought for, the needed number of measurements and the recovery complexity are essentially lowest possible. Moreover, the different vectors $\bm{g}$ that produce the same measurements $\bm{y}$ all have same magnitudes of elements on same positions, meaning that only the phases of the elements differ.
@@
Conversely, 2.2.b. states that when $\bm{x}$ is composed of $s$ frequencies with at least one being non-harmonic, where $n\geq 4s-1$ and the number of measurements $m\geq 8s-3$, there are only two possible solutions (one of them is of course the true vector $\bm{x}$) and they can be recovered with $O(\mathrm{poly}(s))$ complexity. Thus, Fourier phase retrieval is in principle solvable with $O(s)$ measurements and $O(\mathrm{poly}(s))$ recovery complexity for signals containing non-harmonic frequencies and oversampled with at least an oversampling factor of 4. Obviously, to find which of the two possible solutions is $\bm{x}$, one can take an additional measurement with an arbitrary measurement vector for resolving the ambiguity; this is used next to construct $\bm{A}$ for recovering $s$-sparse $\bm{x}$ from phase-less measurements.
@@
        When $m=4s-1$, the elements in $\bm{z}$ are $n$th roots of unity and for almost all $\bm{\theta}$, we have exact recovery of the support of $\bm{x}\in\mathbb{C}_{\leq s}^{n}$ and $|\bm{x}|$ (up to a global real scalar) from $\bm{y}$ and $\bm{A}$. There are $2^{s-1}$ different $\bm{x}$ (all with the same value of $|\bm{x}|$) producing the same $\bm{y}$.
@@
        For almost all $\bm{z}$ and $\bm{\theta}$ with $m\geq 8s-3$, we have exact recovery of the support of $\bm{x}\in\mathbb{C}_{\leq s}^{n}$ and $|\bm{x}|$ (up to a global real scalar) from $\bm{y}$ and $\bm{A}$. There are only two different $\bm{x}$ (with the same value of $|\bm{x}|$) producing the same $\bm{y}$ and they are easily obtainable from each other.
@@
        When the elements in $\bm{z}$ are $n$th roots of unity and for almost all $\bm{\theta}$, $m=4s-1$ measurements suffice to recover $\bm{x}$ from $\bm{y}$ and $\bm{A}$ with a recovery complexity of $O(\mathrm{poly}(s)2^{s})$.
@@
        For almost all $\bm{z}$ and $\bm{\theta}$, $m\geq 8s-3$ suffices to recover $\bm{x}$ from $\bm{y}$ and $\bm{A}$ with $O(\mathrm{poly}(s))$ recovery complexity.
Almost all vectors $\bm{a}$ from a set $\mathbb{A}$ equals the set $\mathbb{A}$ up to a null set.
@@
The results in 4. are obtained rather easily from 2. which can be seen from the similarity of the statements. The result in 4.4.b. directly solves open problem 3.5 in [18] and significantly improves upon it. Namely, that problem asks for an explicit construction of $\bm{A}$ with $m=O(s\log(n/s))$ such that solving (4) results in exact recovery of $\bm{x}$. The combination of the deterministic construction $\bm{F}(\bm{z})^{T}\bm{F}(\bm{\theta})$ and a single measurement as described in 4.4.b. shows that after only $4s$ measurements there is a unique solution to $\bm{x}$ (up to a global phase) in (3); thus even (4) will find it. Noteworthy is that the number of measurements $4s$ in the proposed construction almost achieves the injectivity lower bound of $44s-$ 4. Furthermore, taking $28s-$ 2 measurements instead of $4s$, we are able to reduce the recovery complexity to $O(\mathrm{poly}(s))$ with a novel recovery algorithm, drastically reducing the recovery complexity for large scale problems (large $n$) compared to (4).
@@ II Phase-aware Measurements
  R1: Assume an $n$-dimensional vector $\bm{x}=\bm{V}(\bm{\theta})\bm{g}$, where $\bm{\theta}\in\mathbb{C}^{s}$, $\bm{g}\in\mathbb{C}^{s}$ and $n\geq 2s$. There are samples $\bm{z}\in\mathbb{C}^{m}$, with $m\geq 2s$, such that $\bm{\theta}$ and $\bm{g}$ (and thus $\bm{x}$) can be exactly recovered from
@@
  R2: Assume $n\geq 2s$ and $m\geq 2s$. There is an $m\times n$ measurement matrix construction $\bm{A}$ such that we can exactly recover any $\bm{x}\in\mathbb{C}_{\leq s}^{n}$ from $\bm{y}=\bm{A}\bm{x}$ with $O(\mathrm{poly}(s))$ recovery complexity.

### 06 Feb 2024 v2

@@ Abstract
In this work, we present some new results for compressed sensing and phase retrieval. For compressed sensing, it is shown that if the unknown $n$-dimensional vector can be expressed as a linear combination of $s$ unknown Vandermonde vectors (with Fourier vectors as a special case) and the measurement matrix is a Vandermonde matrix, exact recovery of the vector with $2s$ measurements and $O(\mathrm{poly}(s))$ complexity is possible when $n\geq 2s$. Hence, in this case, there is no need for random measurements (with the number of measurements scaled by $\log(n)$) as required by conventional compressed sensing theory. From these results, a measurement matrix is constructed from which it is possible to recover $s$-sparse $n$-dimensional vectors for $n\geq 2s$ with as few as $2s$ measurements and with a recovery algorithm of $O(\mathrm{poly}(s))$ complexity. This seems to be the first measurement matrix construction achieving exact recovery of $s$-sparse vectors with $O(s)$ measurements (achieving the lower bound of $2s$) and $O(\mathrm{poly}(s))$ complexity (typically, the recovery complexity is $O(\mathrm{poly}(n))$), significantly improving upon a classical construction by Bourgain et al. In the second part of the work, these results are extended to the challenging problem of phase retrieval (compressed sensing with magnitude-only observations). The most significant discovery in this direction is that if the unknown $n$-dimensional vector is composed of $s$ frequencies with at least one being non-harmonic, $n\geq 4s-1$ and we take at least $38s-$ 3 Fourier measurements, there are, remarkably, only two possible vectors producing the observed measurement values and they are easily obtainable from each other. The two vectors can be found by an algorithm with only $O(\mathrm{poly}(s))$ complexity. This result is interesting when put in relation to recent work by Xu which shows that solving Fourier phase retrieval with $O(\mathrm{poly}(n))$ complexity proves that P = NP holds for the well known P vs NP problem. An immediate application of the new result is construction of a measurement matrix from which it is possible to recover almost all $s$-sparse $n$-dimensional signals (up to a global phase) from $O(s)$ magnitude-only measurements and $O(\mathrm{poly}(s))$ recovery complexity when $n\geq 4s-1$. As a consequence, for almost all $s$-sparse $n$-dimensional signals, the proposed matrix construction significantly improves upon a result by Xu on finding a measurement matrix construction from which one can (provably) recover $s$-sparse signals with $O(s\log(n/s))$ phase-less measurements and 1-norm minimization.
@@ I Introduction
where $\mathbb{C}^{n}_{\leq s}$ is the space of $n$-dimensional complex vectors with at most $s$ elements being non-zero and $\bm{A}$ is a known $m\times n$ measurement matrix ($m$ is the number of measurements). This problem has a rich history. Seminal works [1, 2, 3] and references therein show that, remarkably, there exist matrices $\bm{A}$ such that $\bm{x}$ can be recovered with far less measurements than $n$. We will refer to exact recovery of $\bm{x}$ as recovery of almost all $\bm{x}$ (i.e. except for possibly some $\bm{x}$ that all belong to a set of measure zero in $\mathbb{C}_{\leq s}^{n}$); another common phrase is recovery with probability 1. By probabilistic recovery we refer to recovery of $\bm{x}$ at some probability strictly smaller than 1. The obvious 0-norm minimization that searches across all subsets of $s$ columns from $\bm{A}$ and inverts them to obtain a candidate solution is of exponential complexity. Fortunately, the work in [4] shows that there exists a recovery algorithm for (1) with $O(\mathrm{poly}(n))$ complexity, where $O(.)$ stands for big O notation. More specifically, [4] shows that, for a given $\bm{y}=\bm{A}\bm{x}$, the unique solution to the convex 1-norm optimization problem
@@
In many practical instances of compressed sensing and phase retrieval, $\bm{A}$ is constrained to be a Fourier matrix [1, 17, 11] which in practice corresponds to an apparatus that takes frequency measurements of an underlying signal. Compressed sensing and (sparse) phase retrieval with Fourier measurement matrices, known as Fourier compressed sensing and (sparse) Fourier phase retrieval, respectively, are notoriously difficult problems, as described in [1, 6, 17]. The main challenge is in obtaining an affordable recovery complexity (e.g. $O(\mathrm{poly}(n))$) with few measurements (e.g. $O(s\log(n/s))$). The breakthrough work in compressed sensing [1] considered random Fourier measurements in order to escape the deterministic nature of the problem and instead provide probabilistic recovery with complexity $O(\mathrm{poly}(n))$ and $O(s\log(n/s))$ measurements. The contributions in this work on compressed sensing show that if $\bm{x}$ is instead $s$-sparse in a Vandermonde basis, with $n\geq 2s$, then it can be recovered exactly from $O(s)$ deterministic Fourier measurements and $O(\mathrm{poly}(s))$ recovery complexity; hence, there is no need for random measurements in this case. In case of (non-sparse) Fourier phase retrieval, as shown in [6], the reason for why no exact $O(\mathrm{poly}(n))$ complexity algorithms have been proposed so far is because in doing so one solves the well known P vs NP problem, proving that P = NP holds. Interestingly, the contributions in this work on Fourier phase retrieval show that if the signal is $s$-sparse in a non-DFT basis, $n\geq 4s-1$, and we take $O(s)$ measurements, then there are in general two possible signals that give rise to the observed measurements. Moreover, the two signals are connected with a simple transformation, so that knowing one of them makes it easy to compute the other.
@@ I-A Contributions
Let $\mathbb{T}=\{z;z\in\mathbb{C},|z|=1\}$ be the set of all points on the complex unit circle and (.$(.)^{T}$ the matrix transpose operator. The term almost all $\bm{x}\in\mathbb{A}$, for some set $\mathbb{A}$, is referring to a set that equals $\mathbb{A}$ up to a null set (set of measure zero in $\mathbb{A}$). When $z_{k}\in\mathbb{T}$, $10\leq k\leq m-$ 1, the Vandermonde matrix becomes a Fourier matrix, which we for clarity henceforth denote as $\bm{F}(\bm{z})$ (with DFT matrix as a special case when $z_{k}$ are $n$th roots of unity and $m=n$). For Vandermonde and Fourier measurement matrices, we have the following main contributions
@@
      For almost all $\bm{z}\in\mathbb{C}^{m}$, $m\geq 3s$ suffices.
@@
      If the elements in $\bm{\theta}$ are $n$th roots of unity ($\bm{F}(\bm{\theta})$ equals a subset of $s$ columns from a $n\times n$ DFT matrix), we have exact recovery of $\bm{\theta}$ and $|\bm{g}|$ (up to a global real positive scalar) from $\bm{y}$ and $\bm{A}$ with $O(\mathrm{poly}(s))$ recovery complexity. There are $2^{s-1}$ different $\bm{g}$ (up to a global phase), all having the same value of $|\bm{g}|$, producing the same $\bm{y}$.
@@
      If not all elements in $\bm{\theta}$ are $n$th roots of unity and for almost all $\bm{g}$
@@
        When the elements in $\bm{z}$ are $n$th roots of unity and $m=4s-1$, we have exact recovery of $\bm{\theta}$ and $|\bm{g}|$ (up to a global real positive scalar) from $\bm{y}$ and $\bm{A}$ with $O(\mathrm{poly}(s))$ recovery complexity. There are $2^{s-1}$ different $\bm{g}$ (up to a global phase), all having the same value of $|\bm{g}|$, producing the same $\bm{y}$.
@@
        For almost all $\bm{z}\in\mathbb{T}^{m}$ with $m\geq 8s-3$, we have exact recovery of $\bm{\theta}$ and $|\bm{g}|$ (up to a global real positive scalar) from $\bm{y}$ and $\bm{A}$ with $O(\mathrm{poly}(s))$ recovery complexity. There are only two different $\bm{g}$ (up to a global phase), both having the same value of $|\bm{g}|$, that produce the same $\bm{y}$; moreover, the two solutions are easily obtainable from each other.
@@
For phase-less measurements, the result in 2. states that $\bm{\theta}$ and $c|\bm{g}|$, for unknown real positive scalar $c$, can be found with $O(s)$ measurements and $O(\mathrm{poly}(s))$ recovery complexity. Thus, in scenarios where only the frequency components $\bm{\theta}$ and their relative magnitudes are sought for, the needed number of measurements and the recovery complexity are essentially lowest possible. Moreover, the different vectors $\bm{g}$ that produce the same measurements $\bm{y}$ all have same magnitudes of elements on same positions, meaning that only the phases of the elements differ.
@@
Conversely, 2.2.b. states that when $\bm{x}$ is composed of $s$ frequencies with at least one being non-harmonic, where $n\geq 4s-1$ and the number of measurements $m\geq 8s-3$, there are only two possible solutions (one of them is of course the true vector $\bm{x}$) and they can be found with $O(\mathrm{poly}(s))$ complexity. Thus, Fourier phase retrieval is in principle solvable with $O(s)$ measurements and $O(\mathrm{poly}(s))$ recovery complexity for signals containing non-harmonic frequencies and oversampled with at least an oversampling factor of 4. Obviously, to find which of the two possible solutions is $\bm{x}$, one can use additional side information as explained in [6] (e.g. knowing one element from $\bm{x}$) or take an additional measurement with an arbitrary/random measurement vector for resolving the ambiguity; the latter approach is used next to construct $\bm{A}$ for recovering almost all $s$-sparse $\bm{x}$ from phase-less measurements.
@@
        When $m=4s-1$, the elements in $\bm{z}$ are $n$th roots of unity and for almost all $\bm{\theta}$, we have exact recovery of the support of $\bm{x}\in\mathbb{C}_{\leq s}^{n}$ and $|\bm{x}|$ (up to a global real positive scalar) from $\bm{y}$ and $\bm{A}$. There are $2^{s-1}$ different $\bm{x}$ (all with the same value of $|\bm{x}|$) producing the same $\bm{y}$.
@@
        For almost all $\bm{z}$ and $\bm{\theta}$ with $m\geq 8s-3$, we have exact recovery of the support of $\bm{x}\in\mathbb{C}_{\leq s}^{n}$ and $|\bm{x}|$ (up to a global real positive scalar) from $\bm{y}$ and $\bm{A}$. There are only two different $\bm{x}$ (with the same value of $|\bm{x}|$) producing the same $\bm{y}$ and they are easily obtainable from each other.
@@
        When the elements in $\bm{z}$ are $n$th roots of unity and for almost all $\bm{\theta}$, $m=4s-1$ measurements suffice to exactly recover $\bm{x}\in\mathbb{C}_{\leq s}^{n}$ from $\bm{y}$ and $\bm{A}$ with a recovery complexity of $O(\mathrm{poly}(s)2^{s})$.
@@
        For almost all $\bm{z}$ and $\bm{\theta}$, $m\geq 8s-3$ suffices to exactly recover $\bm{x}\in\mathbb{C}_{\leq s}^{n}$ from $\bm{y}$ and $\bm{A}$ with $O(\mathrm{poly}(s))$ recovery complexity.
@@
The results in 4. are obtained rather easily from 2. which can be seen from the similarity of the statements. The result in 4.4.b. solves problem 3.5 in [18] for almost all $\bm{x}\in\mathbb{C}_{\leq s}^{n}$ and significantly improves upon it. Namely, that problem asks for an explicit construction of $\bm{A}$ with $m=O(s\log(n/s))$ such that solving (4) results in recovery of all $\bm{x}\in\mathbb{C}_{\leq s}^{n}$. The combination of the deterministic construction $\bm{F}(\bm{z})^{T}\bm{F}(\bm{\theta})$ and a single arbitrary measurement as described in 4.4.b. shows that after only $4s$ measurements there is a unique solution to $\bm{x}$ (up to a global phase) in (3) for almost all $\bm{x}\in\mathbb{C}_{\leq s}^{n}$; thus even (4) will find it. Noteworthy is that the number of measurements $4s$ in the proposed construction almost achieves the injectivity lower bound of $44s-$ 4. Furthermore, taking $28s-$ 2 measurements instead of $4s$, we are able to reduce the recovery complexity to $O(\mathrm{poly}(s))$ with a novel recovery algorithm, drastically reducing the recovery complexity for large scale problems (large $n$) compared to (4).
@@ II Phase-aware Measurements
  R1: Assume an $n$-dimensional vector $\bm{x}=\bm{V}(\bm{\theta})\bm{g}$, where $\bm{\theta}\in\mathbb{C}^{s}$, $\bm{g}\in\mathbb{C}^{s}$ and $n\geq 2s$. There are samples $\bm{z}\in\mathbb{C}^{m}$, with $m\geq 2s$, such that almost all $\bm{\theta}$ and $\bm{g}$ (and thus almost all $\bm{x}$) can be exactly recovered from
@@
  R2: Assume $n\geq 2s$ and $m\geq 2s$. There is an $m\times n$ measurement matrix construction $\bm{A}$ such that we can recover any $\bm{x}\in\mathbb{C}_{\leq s}^{n}$ from $\bm{y}=\bm{A}\bm{x}$ with $O(\mathrm{poly}(s))$ recovery complexity.